OpenStudy (anonymous):

help me please!!! 775(1 + 0.0425)^1 * 12

OpenStudy (mertsj):

Do you have a calculator?

OpenStudy (anonymous):

use pemdas 775(1.0425)^1*12 775(1.0425)*12 807.9375*12 9695.25

OpenStudy (anonymous):

oh thank you!!! but that is not the answer in my book!!

OpenStudy (anonymous):

oh SORRY.. what do they give you?

OpenStudy (mathmale):

I'm concerned about that 1 near the end of your expression: 775(1 + 0.0425)^1 * 12 Where did that come from?

OpenStudy (mathmale):

Would you mind double checking to ensure that you've copied the problem down here exactly as it appears in your book? Could you use the Equation Editor for greater clarity?

OpenStudy (anonymous):

ok the problem is actually this: $775 deposit earning 4.25% compounded anually, after 12 years

OpenStudy (anonymous):

oh that explains alot...

OpenStudy (anonymous):

:D sorry!!

OpenStudy (anonymous):

its cool. dont worry

OpenStudy (anonymous):

thank you guys!!!

OpenStudy (mathmale):

775(1 + 0.0425) is fine, but I'm not sure I understand your ^1 * 12 at the end. OK, thanks very much for the clarification; that was important! the relevant formula, a simplified version that takes into account the fact that interest is compounded once per year, is A=P (1+r)^t, where P=principal amount, r=interest rate as a decimal fraction, and t=number of years.

OpenStudy (mathmale):

So, JustinDrewBieber, you'll end up with A=775(1+0.0425)^12.

OpenStudy (mathmale):

If you insist on typing in that " 1 " to indicate annual compounding, then please enclose it and the 12 in parentheses: ^(1*12). Otherwise your computer calculates things in the wrong order!

OpenStudy (anonymous):

oh ok thank you!!! but how do I work it now?

OpenStudy (mathmale):

the correct answer is $1,277.07. share half of that with me, please?

OpenStudy (mathmale):

Have a TI-84 calculator?

OpenStudy (mathmale):

Unfortunately, I'm much less familiar with other calcs., but if you have a TI-84, it'd be a snap to walk you thru this calculation.

OpenStudy (mathmale):

JDB?

OpenStudy (anonymous):

IDK!!

OpenStudy (anonymous):

I have TI-30XS

OpenStudy (anonymous):

nevermind, I think I got it!!!

OpenStudy (mathmale):

Don't have one and am less fam. with that. But anyway. I'd suggest you type in 1.0450^12 first. Does that seem workable to you (know how to do that)?

OpenStudy (mathmale):

Then multiply the result by 775 and you've got it. Congrats.

OpenStudy (anonymous):

Thanks SOOOOOOO much!!!!

OpenStudy (mathmale):

My great pleasure. Keep up the good work! And thank you for the medal!